I am a 2 digit prime number
The sum of my digits is divisible by 4
The product of my digits is less then 5
I am greater then 20
What number am I?
What if I am less then 20?
I've been stuck on this question for ages, please help!!
Can you talk me through the steps you take to get the answer?

Answer by Edwin McCravy(20056)

It's good to be able to recognize all the
prime numbers less than 100:

2, 3, 5, 7, 11, 13, 17, 19, 23, 29, 
31, 37, 41, 43, 47, 53, 59, 61, 67,
71, 73, 79, 83, 89, 97.

Those greater than 20 are

23, 29, 31, 37, 41, 43, 47, 53, 59,
61, 67, 71, 73, 79, 83, 89, 97.

The only one that has sum of digits 4 is 31, since 3+1=4.

31 is greater than 20.

The product of its digits is 3×1 = 3, which is less than 5

----------------------------------------------------------

All the two-digit prime numbers less than 20 are

11, 13, 17, 19

The only one with sum of digits 4 is 13, since 1+3=4. 

Its product of digits is 1×3 = 3, which is less than 5.

Edwin